OpenStudy (anonymous):

find the intercepts f(x)=x^2-6-x

OpenStudy (anonymous):

x=-2,3, y=-6

OpenStudy (anonymous):

can you explain how you got this

OpenStudy (anonymous):

the x intercepts are when y = 0 so x^2-6-x = 0 which can be factored; (x-3)(x+2) = 0 so the intercepts are at 3 and -2 the y intercept occurs when x = 0 so plugging x=0 in the expression: y = 0 - 0 - 6 = -6

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

yes can you help me find the intercepts of f(x)=2x^2-3x+5

OpenStudy (anonymous):

ok first do you know how to factor f(x) ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok well this requires some practice but if you do that you'll get it this type of expression is called a quadratic - there is 1 ^x , one term in x and one number first the 2x^2 = 2x multiplied by x so we can write (2x )(x ) now we want the other two numbers: these two numbers are +1 and -5 because you want to get -3x in the middle and 2x1 and 1x -5 gives this so we have (2x - 5)(x + 1) are our factors this is hard to grasp at first for most people . you can find details on the wolframalpha website where it is explained very well

OpenStudy (anonymous):

how about f(x)=3x^2+2x-5

OpenStudy (anonymous):

so back to the problem 2x - 5)(x + 1) = 0 which gives 2x-5 = 0 and so x = 2.5 and x+1 = 0 so x = -1 in tercepts on x axis are -1,2-5 y intersept is when x = 0 this is y = 0 + 0 + 5 = 5

OpenStudy (anonymous):

sorry -1, 2.5 and 5

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

can you answer my next question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

evaluate \[f(x)=\sin^2x+\cos^2x, x=9\pi/4\]

OpenStudy (anonymous):

well sin^2x+cos^2x is always = 1 whatever value x is given so there is no need to evaluate it

OpenStudy (anonymous):

that is a crafty question

OpenStudy (anonymous):

why

OpenStudy (anonymous):

well as i said that f(x) is always = 1, if this is written homework just plug in x=9pi/4 into the expression and write '= 1'.

OpenStudy (anonymous):

and you'll get full credit for it

OpenStudy (anonymous):

ok can you help me with \[(2/\cos \theta)+(4/1-\sin \theta)\]

OpenStudy (anonymous):

what do u want to do with this - simplify it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you have an answer for this

OpenStudy (anonymous):

hmmm (2/cosθ)+(4/1−sinθ) = 2 - 2sinθ + 4cosθ / cosθ(1-sinθ) thats what i've got so far but i 'm struggling a bit - hold on a bit

OpenStudy (anonymous):

ok thank you for your help

OpenStudy (anonymous):

thats ok - try posting that again - i might have missed the way to do this

OpenStudy (anonymous):

\[(2/\cos \theta)+(4/1-\sin \theta)\]

OpenStudy (anonymous):

is this good

OpenStudy (anonymous):

can you plz help me no one else seems to know the answer to this question